Oxford Company has limited funds available for investment and must ration the funds among four competing projects. Selected information on the four projects follows:

Project Investment
Required
Net
Present
Value
Life of
the
Project
(years)
Internal
Rate
of Return
A $ 880,000 $ 441,342 7 24 %
B $ 685,000 $ 366,397 12 20 %
C $ 580,000 $ 224,939 7 21 %
D $ 780,000 $ 169,812 3 22 %

The net present values above have been computed using a 10% discount rate. The company wants your assistance in determining which project to accept first, second, and so forth.

Required:

1. Compute the project profitability index for each project.

2. In order of preference, rank the four projects in terms of net present value, project profitability index and internal rate of return.
Answer #1
1
Project profitability index = Net Present Value/Investment Required
Project Project profitability index
A 0.50 =441342/880000
B 0.53 =366397/685000
C 0.39 =224939/580000
D 0.22 =169812/780000
2
Net Present Value Project profitability index Internal Rate of Return
First preference A B A
Second preference B A D
Third preference C C C
Fourth preference D D B